Developing a deep understanding of linear relationships is an important part of the algebra curriculum. One way to analyze these relationships is to examine the way linear equations are represented symbolically, before they are graphed. Review the Semantic Feature Analysis activity, and then think about the following:
Do the graphs of y = x and y = x + 2 intersect with each other?

How could you answer this question without actually graphing each equation?

How would you describe the relationship between the symbolic representation for these lines and their graphs?

Answer by stanbon(75887)

Developing a deep understanding of linear relationships is an important part of the algebra curriculum. One way to analyze these relationships is to examine the way linear equations are represented symbolically, before they are graphed. Review the Semantic Feature Analysis activity, and then think about the following:
Do the graphs of y = x and y = x + 2 intersect with each other?
No because they have the same slope but different y-intercepts.
So they are parallel lines
-------------------------------------------
How could you answer this question without actually graphing each equation?
How would you describe the relationship between the symbolic representation for these lines and their graphs?
y = x+2 is "2" above y = x for every value of x.
=======================================
Cheers,
Stan H.